OpenStudy (anonymous):

Find the product: (3x + 4)(3x - 4)

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\]

OpenStudy (anonymous):

9x^2-12x+12x-16 9x^2-16

OpenStudy (y2o2):

(3x)²-4² = 9x²-16